How do you draw a spinner that fits all four clues:
1. The probability of spinning a 2 is equal to the probability of spinning a 1.
2. The probability of spinning all the numbers is not equal.
3. The probability of spinning a number other than a 1 is 75%
4. There are three numbers on the spinner.
I can't show how I draw but I can illustrate it. In the spinner, the numbers are 1,1,2,3
I have P(1) = 2/4 or 1/2, P(2)= 2/4 = 1/2, P(3) = 3/4 or 75%
Can You tell me did I get it right? If not, how do you solve this problem? Thanks again.

Answer by sudhanshu_kmr(1152)


by using four clues we get:
number of 2 and number of 1 is equal. i.e x
number of other numbers except 1 is 3x
(because probability of 1 is 25% only)
so, number of 1 = x

number of 2 = x

number of third number i.e 3 = 2x
there are total: 4x
P(1) = 1/4
P(2) = 1/4
P(3) = 2/4 = 1/2

IF ANY DOUBT, PLEASE CONTACT..